Please make sure your answers to the following questions are in Radical Form.
Grace and Chandler met for lunch. At 1 p.m., they parted ways. Maybe forever, considering how they left things. Chandler drove due south at 30 mph and Grace drove due east at 60 mph. Apparently, she was more upset than he was. At 2:30 p.m., how far away are Grace and Chandler?
It says that I have to use the Pythagorean Theorem to write an equation that will help solve my problem. Then I must write in a sentence my answer, and explaining what it means. I understand what that is and how you do it. I just don't understand how to do it all in Radical Form.

Answer by stanbon(75887)

Grace and Chandler met for lunch. At 1 p.m., they parted ways. Maybe forever, considering how they left things. Chandler drove due south at 30 mph and Grace drove due east at 60 mph. Apparently, she was more upset than he was. At 2:30 p.m., how far away are Grace and Chandler?
-----
Chandler distance:: 30*(3/2 hr) = 45 miles
Grace distance:: 60(3/2 hr) = 90 miles
-----
Draw the right triangle with legs of 45 and 90.
Ans: dist = hypotenuse = sqrt[45^2+90^2] = sqrt(10125)
= sqrt[25*81*5] = 5*9sqrt(5) = 45sqrt(5) miles
--------
Cheers,
Stan H.